# CinePile: A Long Video Question Answering Dataset and Benchmark

### v1

@@ Appendix F Evaluation on Captioning Metrics
mPLUG-Owl [40] 0.46 1.36 0.31
Video-ChatGPT [17] 0.48 1.14 0.32
MovieChat [26] 0.40 0.66 0.18
@@
Therefore, in this section, we compute traditional video-caption evaluation metrics that emphasize the semantic similarity between the answer key text and the raw model response, instead of exact string matching. We focus our evaluation and discussion on open-source models here, as we qualitatively noted that proprietary models, such as GPT-4V, Gemini-Pro, and Claude, strictly adhere to the prompt instructions, producing only the option letter in their response. Specifically, we calculate the following video-captioning metrics – BERTScore [45], CIDEr [31], and ROUGE-L [15]. BERTScore calculates the contextual similarity between the answer key and model response in the embedding space of a pretrained transformer model like BERT-Base. Calculating the similarity between the latent representations, instead of direct string matching, provides robustness to paraphrasing differences in the answer key and model response. In contrast, CIDEr evaluates the degree to which the model response aligns with the consensus of a set of reference answer keys. In our setup, each question is associated with only one reference answer. The alignment here is computed by measuring the similarity between the non-trivial n-grams present in the model response and the answer key. Finally, ROUGE-L computes the similarity between the answer key and model response based on their longest common subsequence.
We evaluate three open source models, i.e. mPLUG-Owl, Video-ChatGPT, and MovieChat using the aforementioned metrics and report the results in Table 7. In line with the accuracy trend in the main paper, we observe that while mPLUG-Owl and Video-ChatGPT don’t have a substantial difference in their performance, they both significantly outperform MovieChat. These findings further support the reliability of our normalization and post-processing steps during accuracy computation, and underscore the need for improving open-source video-LLMs to close the gap with proprietary models.
Appendix G Example Degenerate Questions
@@
Appendix H Human Study Details

### v3

@@ Appendix H Additional Evaluation Strategies
This is the base.”
What action does Darren take after observing the broken sensor?
- A) Calls for help
- B) Repairs it
- C) Panics
- D) Ignores it
- E) Suggests next steps
[…]
Where is Jelena in relation to the helicopter at the end of the scene?
- A) Inside
- B) In front
- C) To the side
- D) Directly underneath
- E) On it
How do Darren and Reed first interact with each other in the scene?
- A) They argue
- B) They fight
- C) They exchange information
- D) They greet each other
- E) They ignore each other
Table 8: Performance of various models on CinePile ’s test split, as evaluated using various video captioning metrics – BERTSCoRE (Devlin et al., 2018), CIDEr (Vedantam et al., 2015), ROUGE-L (Lin, 2004).
[…]
mPLUG-Owl Ye et al. (2023a) 0.38 0.74 0.22
Video-ChatGPT Maaz et al. (2023) 0.39 0.63 0.23
Intern-VL-2 (1B)  Song et al. (2023) 0.40 1.33 0.28
CogVLM-2  Song et al. (2023) 0.45 1.20 0.31
@@
Therefore, in this section, we compute traditional video-caption evaluation metrics that emphasize the semantic similarity between the answer key text and the raw model response, instead of exact string matching. We focus our evaluation and discussion on open-source models here, as we qualitatively noted that proprietary models, such as GPT-4V, Gemini-Pro, and Claude, strictly adhere to the prompt instructions, producing only the option letter in their response. Specifically, we calculate the following video-captioning metrics – BERTScore (Zhang et al., 2019), CIDEr (Vedantam et al., 2015), and ROUGE-L (Lin, 2004). BERTScore calculates the contextual similarity between the answer key and model response in the embedding space of a pretrained transformer model like BERT-Base. Calculating the similarity between the latent representations, instead of direct string matching, provides robustness to paraphrasing differences in the answer key and model response. In contrast, CIDEr evaluates the degree to which the model response aligns with the consensus of a set of reference answer keys. In our setup, each question is associated with only one reference answer. The alignment here is computed by measuring the similarity between the non-trivial n-grams present in the model response and the answer key. Finally, ROUGE-L computes the similarity between the answer key and model response based on their longest common subsequence.
We evaluate four open source models, i.e. mPLUG-Owl, Video-ChatGPT, Intern-VL-2 (1B), and CogVLM2, using the aforementioned metrics and report the results in Table 8. In line with the accuracy trend in the main paper. These findings further support the reliability of our normalization and post-processing steps during accuracy computation.
Appendix I Human Study Details
Post the study, we interviewed each participant after the survey to ask if they found any systematic issues in any of the questions they were asked to answer about the video. Later, a panel of authors audited all questions where humans got the answer wrong. We noticed that most of the time when a human got a question wrong it was likely due to one of the following reasons (i) due to their inability to attend over the entire clip at once, (ii) due to their inability to understand the dialogue or understand cultural references (iii) carelessness in answering, as the correct answer was indeed present in the video. We did notice some problematic patterns with a small subset of questions. The main issue is distractor similarity, where humans found two plausible answers and they chose one randomly. We present a few such examples in Fig. 11. We removed the questions from the test set for which we found ambiguous answers.
We again conducted a second human study on the test set’s final version, and the human accuracy is 73%. The authors have independently taken the survey, and the corresponding accuracy is 86%. Once again, a careful investigation by a team of authors indicates that even most of these wrong answers are due to human error and confusion over the many events in a scene. We conclude from this study that many of the questions are answerable but difficult. We present the question category-level performance in Sec. 4 in the main paper.
Appendix J Example Degenerate Questions
@@ Appendix I Human Study Details
Appendix K Additional Evaluation Results
K.1 Frame Rate Ablation
In this section we perform an ablation to investigate the utility of visual frames (from a model’s perspective) by completely remove the visual frames and experiment solely with the provided dialogue when evaluating Video-LLMs. We do exactly this in Table 10, and observe that for all models, except Video-ChatGPT, performance significantly declines when evaluated with "only subtitles." This effect is more pronounced in commercial models compared to open-source ones. It appears that better overall models also tend to utilize visual information more effectively.
Table 10: Performance of models with video and subtitles (base case), and when only with subtitles on a subset of CinePile. TEMP - Temporal, CRD - Character and Relationship Dynamics, NPA - Narrative and Plot Analysis, STA - Setting and Technical Analysis, TH - Thematic Exploration.
Model Average CRD NPA STA TEMP TH
Gemini 1.5 Pro 51.72 51.61 56.25 55.45 40.62 50.00
(Only Subtitles) Gemini 1.5 Pro 34.53 35.87 44.44 31.35 32.60 36.36
GPT-4o 50.45 51.14 66.66 52.54 34.78 45.45
(Only Subtitles) GPT-4o 37.23 45.03 44.44 29.66 28.26 45.45
Video-LLaMA2 38.44 45.80 40.74 36.44 19.56 54.54
(Only Subtitles) Video-LLaMA2 33.33 41.22 40.74 27.11 17.39 45.45
Video-ChatGPT 12.92 16.80 3.70 12.82 6.52 20.00
(Only Subtitles) Video-ChatGPT 16.16 22.04 11.53 12.71 13.04 9.09
K.2 Performance on Hard-Split
Appendix L QA Generation Prompt
As the curator of an advanced cinema analysis quiz, your expertise lies in designing intricate and diverse multiple-choice questions with corresponding answers that span the entire spectrum of film analysis.
- Objective: Create diverse and challenging questions based on the film analysis spectrum templates provided below. This spectrum is divided into five subcategories, each comprising several templates. Each template includes a title and a corresponding prototypical question or guideline. Avoid directly replicating the template title and these prototypical questions. Instead, your questions should reflect these elements’ essence, even if not explicitly using the category titles in the question’s wording.
Mandatory Guidelines:
- Template Use: Use the provided question templates as a strict guide, ensuring that your questions are both relevant to the scene and varied in their analytical perspective. The prototype question in each template is for inspiration and should not be copied. Your questions should subtly reflect the prototype’s essence, tailored to the specifics of the scene.
- Sub-Category Balance: Ensure to generate an equal number of questions from each subcategory. This balance is crucial to cover a wide range of analytical perspectives.
- Question and Answer Format:
- Selected Template: Indicate the film analysis Sub-Category and corresponding template your question is inspired by, without restricting the question’s phrasing to the template’s title.
- Questions: Limited to one or two lines, formulated to be insightful and not overtly indicative of the answer. Avoid using direct template titles or overly descriptive language that could hint at the correct answer.
- Answers: Five options per question, formatted as "- A), - B), - C), - D), and - E)", concise and reflective of the question’s depth.
- Answer Key: Specify the correct answer clearly with the formatting, "Correct Answer:", in the line following all the answer options.
- Rationale: Write a rationale explaining the correctness of the "Answer Key" based on the scene’s context in the next line.
Input Information Format:
- Movie scene details will be provided in a structured format comprising two distinct categories, and the relevant scene description. The two categories are as follows:
- <subtitle> for character dialogues (to be used only for identifying character presence, not actions or dialogue content).
- <visual descriptions> for noting characters’ presence, attributes, thematic elements, etc., within the scene.
Movie Scene: {MOVIE_SCENE_TS}
- Spectrum of Film Analysis with Templates:
Sub-Category: Character Analysis
{TEMPLATES_CHAR}
Sub-Category: Narrative Understanding
{TEMPLATES_NARV}
Sub-Category: Scene Setting
{TEMPLATES_SETTING}
Sub-Category: Temporal
{TEMPLATES_TEMPORAL}
Sub-Category: Theme
{TEMPLATES_THEME}
Instructions: Your task is to generate clear, unique, and insightful question-answer pairs strictly following the provided templates. Ensure the distribution of questions covers all subcategories evenly. Strictly avoid using words in the questions that give a strong hint about the answer. You can achieve this by keeping the questions concise and not using too many adjectives or adverbs in the question. Incorrect answers must be plausible and closely mirror the correct answer in length and form. The correct answer should not be deducible solely from the question and/or the wrong answers. After presenting all the options, the correct answer must be distinctly specified, but separate from the list of choices. Additionally, provide a concise rationale about why the question-answer falls into one of the selected templates from the Spectrum of Film Analysis by giving verbatim evidence from the subtitles and/or visual descriptions in the movie scene information.